OpenStudy (anonymous):

Please Help!! Find an equation in standard form for the ellipse with the vertical major axis of length 6, and minor axis of length 4.

OpenStudy (anonymous):

\[\frac{ x ^{2} }{ 2 }+\frac{ y^{2} }{ 3 }=1 \] \[\frac{ x ^{2} }{ 9 } +\frac{ y ^{2} }{ 4 }=1\] \[\frac{ x ^{2} }{ 3 }+\frac{ y ^{2} }{ 2 }=1\] \[\frac{ x ^{4} }{4 }+\frac{ y ^{2} }{ 9 }=1\]

OpenStudy (anonymous):

for the last one it should be x^2

OpenStudy (anonymous):

@SolomonZelman @precal

OpenStudy (anonymous):

@mathstudent55

OpenStudy (solomonzelman):

I haven't done ellipses for a while... I'll read my book to rvw this for a little and if no one helps....

OpenStudy (anonymous):

ok, thank you :)

OpenStudy (solomonzelman):

standard form of the ellipse is assuming that the center is the origin ?

OpenStudy (solomonzelman):

Assuming that (0,0) (the origin) is the center  \(\LARGE\color{blue}{ \frac{x^2}{2^2} +\frac{y^2}{3^2}\rm =1 }\) \(\LARGE\color{black}{ \frac{x^2}{4} +\frac{y^2}{9}=1~~~~~~~\frak {✓} }\)

OpenStudy (solomonzelman):

saying that vertical axis is 6 so 3 from the (0,0) up, and 3 down. An so we go 2 and 2 horizontally

OpenStudy (anonymous):

oh my gosh i get it!

OpenStudy (anonymous):

thank you!!

OpenStudy (solomonzelman):

Anytime !

OpenStudy (anonymous):

can you check if i got this one right?

OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation \[\frac{ x ^{2} }{ 100}+\frac{ y ^{2} }{ 64}=100\]

OpenStudy (anonymous):

i said: Center: (0, 0); Vertices: (-10, 0), (10, 0); Foci: (-6, 0), (6, 0)

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Looks like you are correct.

OpenStudy (anonymous):

oh geez, that second 100 should be a 1

OpenStudy (anonymous):

ok, thank you :)

OpenStudy (solomonzelman):

did you just now say that the question is wrong ?

OpenStudy (anonymous):

no the question is right, i wrote it wrong, it should be \[\frac{ x ^{2} }{ 100}+\frac{ y^{2} }{ 64}=1\]

OpenStudy (solomonzelman):

Oh.. that! I realized that :) Or otherwise I would have told you to divide everything by 100 .

OpenStudy (anonymous):

ohh haha ok :)